# Conformal Prediction Regions are Imprecise Highest Density Regions

### v1

@@ Abstract.
Recently, Cella and Martin proved how, under an assumption called consonance, a credal set (i.e. a closed and convex set of probabilities) can be derived from the conformal transducer associated with transductive conformal prediction. We show that the Imprecise Highest Density Region (IHDR) associated with such a credal set corresponds to the classical Conformal Prediction Region. In proving this result, we relate the set of probability density/mass functions (pdf/pmf’s) associated with the elements of the credal set to the imprecise probabilistic concept of a cloud. As a result, we establish new relationships between Conformal Prediction and Imprecise Probability (IP) theories. A byproduct of our presentation is the discovery that consonant plausibility functions are monoid homomorphisms, a new algebraic property of an IP tool.
@@ 2.1. Conformal Prediction
Suppose that there is an exchangeable process $Y_{1},Y_{2},\ldots$ with distribution $\mathbf{P}$, where each $Y_{i}$ is a random element taking values in $\mathbb{Y}$. Recall that a sequence is exchangeable if, for any $k\in\mathbb{N}$ and any permutation $\tau$, the two random vectors $(Y_{1},\ldots,Y_{k})^{\top}$ and $(Y_{\tau(1)},\ldots,Y_{\tau(k)})^{\top}$ have the same joint distribution. This implies that the marginal distributions of the $Y_{i}$’s are the same. [^2] Exchangeability does not imply independence, so the standard i.i.d. setup is in a sense more restrictive. Note also that we are not assuming any parametric form for the distribution $\mathbf{P}$.
@@
Transductive conformal prediction (TCP) is not the only way of carrying out conformal prediction. There exists another type called inductive (or split) conformal prediction (ICP). It assumes exchangeability and it is used to build the same region $\mathscr{R}_{\alpha}(\mathbf{y^{n}})$ in (1) having the same guarantee (2), while being computationally less expensive than TCP. Since ICP produces the same region as TCP, which is what we need in light of the goal of our paper, we do not present ICP in detail. We refer the interested reader to Papadopoulos, (2008); Angelopoulos et al., (2024).
@@ 2.2. Imprecise Probabilities
A consonant plausibility function $pl$ is a monoid homomorphism.
@@
Note that $\pi(y)$ is the upper probability assigned to the singleton set $\{y\}$, which means that the entire lower and upper probability pair $(\underline{P},\overline{P})$ is determined by the plausibility assigned to singletons. This is a unique feature of the consonance model (Cella and Martin,, 2022; Dubois and Prade,, 1990). If we denote by $\mathcal{F}_{m}$ the set of all focal elements, i.e., $A\in\Sigma_{\mathbb{Y}}$ such that $m(A)>0$, then consonance corresponds to the case of having nested focal elements. Specifically, for any two elements $A,B\in\Sigma_{\mathbb{Y}}$ we have either $A\subseteq B$ or $B\subseteq A$, indicating a hierarchical structure (Shafer,, 1976).
@@ 3. Relationship between Belief Functions and TCP
One of the main concerns in the conformal literature is to identify function $\pi$, taking values in $[0,1]$, such that $\pi(Y_{n+1},\mathbf{Y^{n}})$ is stochastically no smaller than $\text{Unif}(0,1)$ under any exchangeable $P\in\Delta_{\mathbb{Y}}$, where $\Delta_{\mathbb{Y}}$ denotes the space of all probability measures on $\mathbb{Y}$. Also compare this to classical hypothesis testing in statistics, where the p-values under a true null hypothesis are distributed uniformly. Vovk et al., (2005, Theorem 11.1) show that the conformal transducer returned by Algorithm 1 satisfies this requirement. Suppose now that function $\pi$ also satisfies
@@ 4. Imprecise Highest Density Regions
Like in the case of conformal prediction, the goal is to predict $Y_{n+1}$ using a method that is valid. To this end, an IHDR proceeds from a credal set $\mathcal{P}_{\text{pred}}$ of candidate distributions on $\mathbb{Y}$, which is supposed to be reliable in the sense that is contains the true distribution. [^7] Subscript “pred” denotes the fact that, in the present paper, we are interested in a credal set of predictive distributions derived from the data $\mathbf{y^{n}}$ at hand via some procedure, e.g. generalized Bayes’ or geometric updating (Gong and Meng,, 2021).
@@ 5. Building the Equivalent Prediction Region
Consider the collection
$\mathbf{\mathfrak{q}}\coloneqq\left\{{q:q=\frac{\text{d}Q}{\text{d}\mu}\text{, }Q\in\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})}\right\},$ (10)
where $\mu$ is a sigma-finite dominating measure (e.g. the Lebesgue measure if $\mathbb{Y}$ is uncountable or the counting measure if $\mathbb{Y}$ is at most countable), and call $\underline{\pi}$ its lower envelope, that is, $\underline{\pi}(\cdot)=\inf_{q\in\mathbf{\mathfrak{q}}}q(\cdot)$. Then, by the consonance property of $\pi$, we have that $[\underline{\pi},\pi]$ is a cloud (Augustin et al.,, 2014, Definition 4.6). Neumaier’s probabilistic constraint on clouds (Neumaier,, 2002), (Augustin et al.,, 2014, Equation (4.9)) give us the following
@@
$>P[Y_{n+1}\in\{y\in\mathbb{Y}:\pi(y,\mathbf{y^{n}})>\alpha\}]$
@@
$\begin{split}\underline{\Pi}_{\mathbf{y^{n}}}[Y_{n+1}&\in\mathscr{R}_{\alpha}(\mathbf{y^{n}})]\\ &\geq\underline{\Pi}_{\mathbf{y^{n}}}[Y_{n+1}\in\{y\in\mathbb{Y}:\pi(y,\mathbf{y^{n}})>\alpha\}]\\ &=1-\alpha=\underline{\Pi}_{\mathbf{y^{n}}}(Y_{n+1}\in\text{IR}_{\alpha}^{\mathcal{M}}),\end{split}$ (11)
@@ 5.1. Varying the Nonconformity Measure $\Psi$
Proposition 7 tells us that, while it is true that the CPR $\mathscr{R}_{\alpha}^{\Psi}(\mathbf{y^{n}})$ corresponds to the IHDR $\text{IR}^{\mathcal{M}}_{\alpha}$ associated with the conformally-built credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ for a fixed non-conformity measure $\Psi$, we can always find a new non-conformity measure $\Psi^{\prime}$ that further improves on $\text{IR}^{\mathcal{M}}_{\alpha}=\mathscr{R}_{\alpha}^{\Psi}(\mathbf{y^{n}})$.
@@ 5.2. Discussion
Before concluding this section, let us add a discussion of the results we presented so far. As pointed out by Ryan Martin and coauthors in their recent works – see e.g. (Cella and Martin,, 2023; Martin,, 2023, 2022), and especially (Cella and Martin,, 2021) – there seems to be a sense in which the conformal prediction procedure assumes a vacuous prior, which captures the concept of maximal (prior) ambiguity. That is, conformal prediction appears to be akin to a Bayesian Sensitivity Analysis (BSA) procedure (Berger,, 1984) whose prior credal set is the whole space of parameter probabilities. In BSA, though, the posterior parameter probabilities (and hence, also the predictive probabilities on $\mathbb{Y}$) obtained from this prior class are again vacuous: no finite sample is enough to annihilate a sufficiently extreme prior belief (Pericchi,, 1998), (Caprio et al., 2024a, , Appendix B). On the contrary, conformal prediction is able to derive a non-vacuous predictive credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})\subsetneq\Delta_{\mathbb{Y}}$.
@@
Another issue worth discussing is that the consonance assumption induces a “distortion” in the shape of the credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ induced by the conformal transducer $\pi$. With this, we mean that $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ is “pushed” towards the boundaries of the space of probabilities on $\mathbb{Y}$. To see this, consider the following example.
@@ 6. Conclusion
Finally, we would like to further our investigation on the shortcomings of the consonance assumption. This is because Caprio et al., (2025) showed that there is at least a case of conformalized credal sets for classification problems in which such an assumption can be foregone.

### v2

@@ Abstract.
Recently, Cella and Martin proved how, under an assumption called consonance, a credal set (i.e. a closed and convex set of probabilities) can be derived from the conformal transducer associated with transductive conformal prediction. We show that the Imprecise Highest Density Region (IHDR) associated with such a credal set corresponds to the classical Conformal Prediction Region. In proving this result, we establish a new relationship between Conformal Prediction and Imprecise Probability (IP) theories, via the IP concept of a cloud. A byproduct of our presentation is the discovery that consonant plausibility functions are monoid homomorphisms, a new algebraic property of an IP tool.
@@ 2.1. Conformal Prediction
Suppose that there is an exchangeable process $Y_{1},Y_{2},\ldots$ with distribution $\mathbf{P}$, where each $Y_{i}$ is a random element taking values in $\mathbb{Y}$. Throughout the paper, we denote by $\mathbf{P}$ the true data generating process, and by $P$ a generic element of $\Delta_{\mathbb{Y}}$, the space of all (countably additive) probability measures on $\mathbb{Y}$. Recall that a sequence is exchangeable if, for any $k\in\mathbb{N}$ and any permutation $\tau$, the two random vectors $(Y_{1},\ldots,Y_{k})^{\top}$ and $(Y_{\tau(1)},\ldots,Y_{\tau(k)})^{\top}$ have the same joint distribution. This implies that the marginal distributions of the $Y_{i}$’s are the same. [^2] Exchangeability does not imply independence, so the standard i.i.d. setup is in a sense more restrictive. Note also that we are not assuming any parametric form for the distribution $\mathbf{P}$.
@@
Transductive conformal prediction (TCP) is not the only way of carrying out conformal prediction. There exists another type called inductive (or split) conformal prediction (ICP). It assumes exchangeability and it is used to build the same region $\mathscr{R}_{\alpha}(\mathbf{y^{n}})$ in (1) having the same guarantee (2), while being computationally less expensive than TCP. Since ICP produces the same region as TCP, which is what we need in light of the goal of our paper, we do not present ICP in detail. We refer the interested reader to (Papadopoulos,, 2008; Angelopoulos et al.,, 2024).
@@ 2.2. Imprecise Probabilities
A consonant plausibility function $pl$ is a monoid homomorphism between the monoids $(\Sigma_{\mathbb{Y}},\cup)$ and $([0,1],\oplus)$, where $\cup$ is the set union operation and $\oplus$ is the tropical addition on $[0,1]$
@@
While at the moment Lemma 1 seems like a result of independent interest, unrelated to our main result in Section 5, we point out how in the future it may be useful to relate Imprecise Probabilities to an algebraic version of Conformal Prediction, in the spirit of the field of Algebraic Statistics (Sullivant,, 2018).
Note that $\pi(y)$ is the upper probability assigned to the singleton set $\{y\}$, which means that the entire lower and upper probability pair $(\underline{P},\overline{P})$ is determined by the plausibility assigned to singletons. This is a unique feature of the consonance model (Cella and Martin,, 2022; Dubois and Prade,, 1990). If we denote by $\mathcal{F}_{m}$ the set of all focal elements, i.e., $A\in\Sigma_{\mathbb{Y}}$ such that $m(A)>0$, then consonance corresponds to the case of having nested focal elements. Specifically, for any two elements $A,B\in\Sigma_{\mathbb{Y}}$ we have either $A\subseteq B$ or $B\subseteq A$, indicating a hierarchical structure (Shafer,, 1976). We mention in passing that in classical possibility theory, “consonance” refers to the nestedness property we just described, while condition $\sup_{y\in\mathbb{Y}}\pi(y)=1$ is usually referred to as “normalization”. We do not make a distinction between the two to be consistent with the terminology in Cella and Martin’s works, and because normalization combined with the condition $pl(A)=\sup_{y\in A}\pi(y)$, $A\in\Sigma_{\mathbb{Y}}$, implies nestedness.
@@ 3. Relationship between Belief Functions and TCP
One of the main concerns in the conformal literature is to identify function $\pi$, taking values in $[0,1]$, such that $\pi(Y_{n+1},\mathbf{Y^{n}})$ is stochastically no smaller than $\text{Unif}(0,1)$ under any exchangeable $P\in\Delta_{\mathbb{Y}}$. Also compare this to classical hypothesis testing in statistics, where the p-values under a true null hypothesis are distributed uniformly. Vovk et al., (2005, Theorem 11.1) show that the conformal transducer returned by Algorithm 1 satisfies this requirement. Suppose now that function $\pi$ also satisfies
@@ 4. Imprecise Highest Density Regions
Like in the case of conformal prediction, the goal is to predict $Y_{n+1}$ using a method that is valid. To this end, an IHDR proceeds from a credal set $\mathcal{P}_{\text{pred}}$ of candidate distributions on $\mathbb{Y}$, which is supposed to be reliable in the sense that it contains the true distribution. [^7] Subscript “pred” denotes the fact that, in the present paper, we are interested in a credal set of predictive distributions derived from the data $\mathbf{y^{n}}$ at hand via some procedure, e.g. generalized Bayes’ or geometric updating (Gong and Meng,, 2021).
@@ 5. Building the Equivalent Prediction Region
Consider the function $\gamma:\mathbb{Y}\rightarrow[0,1]$,
$y\mapsto\gamma(y)=\begin{cases}\pi(y)&\text{if }\pi(y)\leq 0.5\\ 1-\pi(y)&\text{if }\pi(y)>0.5\end{cases}.$
It is easy to see that $\gamma(y)\leq\pi(y)$, for all $y\in\mathbb{Y}$. In addition, by the consonance property of $\pi$, there exists $\tilde{y}\in\mathbb{Y}$ such that $\gamma(\tilde{y})=0$. In turn, we have that $[\gamma,\pi]$ is a cloud (Augustin et al.,, 2014, Definition 4.6). Neumaier’s probabilistic constraint on clouds (Neumaier,, 2002), (Augustin et al.,, 2014, Equation (4.9)) gives us the following
@@
$=P[Y_{n+1}\in\{y\in\mathbb{Y}:\pi(y,\mathbf{y^{n}})>\alpha\}]$
@@
$\begin{split}\underline{\Pi}_{\mathbf{y^{n}}}[Y_{n+1}&\in\mathscr{R}_{\alpha}(\mathbf{y^{n}})]\\ &=\underline{\Pi}_{\mathbf{y^{n}}}[Y_{n+1}\in\{y\in\mathbb{Y}:\pi(y,\mathbf{y^{n}})>\alpha\}]\\ &\geq 1-\alpha=\underline{\Pi}_{\mathbf{y^{n}}}(Y_{n+1}\in\text{IR}_{\alpha}^{\mathcal{M}}).\end{split}$ (10)
@@
Let us also point out that Proposition 5 is an immediate consequence of Couso et al., (2001, Theorem 2), where it is also shown that Corollary 6.1 holds for all $P\in\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$. Our result generalizes it slightly, by showing that the statement actually holds for all exchangeable $P$ on $\mathbb{Y}$.
@@ 5.1. Varying the Nonconformity Measure $\Psi$
Proposition 7 tells us that, while it is true that the CPR $\mathscr{R}_{\alpha}^{\Psi}(\mathbf{y^{n}})$ corresponds to the IHDR $\text{IR}^{\mathcal{M}}_{\alpha}$ associated with the conformally-built credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ for a fixed non-conformity measure $\Psi$, we can always find a new non-conformity measure $\Psi^{\prime}$ that further improves on $\text{IR}^{\mathcal{M}}_{\alpha}=\mathscr{R}_{\alpha}^{\Psi}(\mathbf{y^{n}})$. In principle, different nonconformity measures may reduce the size of the credal set; in addition, further shrinking the set may eventually require more data.
@@ 5.2. Discussion
Before concluding this section, let us add a discussion of the results we presented so far. As pointed out by Ryan Martin and coauthors in their recent works (see e.g. (Cella and Martin,, 2023; Martin,, 2023, 2022), and especially (Cella and Martin,, 2021)), there seems to be a sense in which the conformal prediction procedure assumes a vacuous prior, which captures the concept of maximal (prior) ambiguity. That is, conformal prediction appears to be akin to a Bayesian Sensitivity Analysis (BSA) procedure (Berger,, 1984) whose prior credal set is the whole space of parameter probabilities. In BSA, though, the posterior parameter probabilities (and hence, also the predictive probabilities on $\mathbb{Y}$) obtained from this prior class are again vacuous: no finite sample is enough to annihilate a sufficiently extreme prior belief (Pericchi,, 1998), (Caprio et al., 2024a, , Appendix B). On the contrary, conformal prediction is able to derive a non-vacuous predictive credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})\subsetneq\Delta_{\mathbb{Y}}$.
@@
Another issue worth discussing is that the consonance assumption induces a “distortion” in the shape of the credal set $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ induced by the conformal transducer $\pi$. With this, we mean that $\mathcal{M}(\overline{\Pi}_{\mathbf{y^{n}}})$ is “pushed” towards the boundaries of the space of probabilities on $\mathbb{Y}$, and that the true data generating process is not its centroid. To see this, consider the following example.
@@
Finally, we point out how proving or disproving a fundamental incompatibility between consonance and minimality (i.e. whether consonance might prevent the credal set from being minimal in size) is a highly non-trivial question, which we plan to inspect in future research.
@@ 6. Conclusion
Finally, we would like to further our investigation on the shortcomings of the consonance assumption. This is because (Caprio et al.,, 2025) showed that there is at least a case of conformalized credal sets for classification problems in which such an assumption can be foregone.
Acknowledgments
Michele Caprio was partially funded by the Army Research Office, grant number ARO MURI W911NF2010080. He wishes to express his gratitude to Ryan Martin and Marianne Johnson for insightful discussions. Yusuf Sale is supported by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. All the authors wish to express their gratitude to four anonymous reviewers for their insightful comments.